# Msmsfnet: a multi-stream and multi-scale fusion net for edge detection

### v1

@@ Abstract
Edge detection is a long standing problem in computer vision. Recent deep learning based algorithms achieve state-of-the-art performance in publicly available datasets. Despite the efficiency of these algorithms, their performance, however, relies heavily on the pretrained weights of the backbone network on the ImageNet dataset. This limits heavily the design space of deep learning based edge detectors. Whenever we want to devise a new model, we have to train this new model on the ImageNet dataset first, and then fine tune the model using the edge detection datasets. The comparison would be unfair otherwise. However, it is usually not feasible for many researchers to train a model on the ImageNet dataset due to the limited computation resources. In this work, we study the performance that can be achieved by state-of-the-art deep learning based edge detectors in publicly available datasets when they are trained from scratch, and devise a new network architecture, the multi-stream and multi-scale fusion net (msmsfnet), for edge detection. We show in our experiments that by training all models from scratch to ensure the fairness of comparison, out model outperforms state-of-the-art deep learning based edge detectors in three publicly available datasets.
Index Terms:
@@
Edge detection, deep learning, pretrained weights, ImageNet, msmsfnet.
publicationid: pubid:
I Introduction
Edge detection in both natural images [1, 2, 3, 4, 5, 6] and remote sensing images [7, 8, 9, 10, 11] has been studied for several decates due to the importance of edge features in image analysis and the use of edges in many applications such as line segment detection [12, 13, 14], image segmentation [5, 15], optical flow estimation [16] and object detection [17, 18]. Despite the huge amount of researches that are dedicated to the problem of edge detection, state-of-the-art edge detection algorithms mainly rely on deep learning. Convolutional Neural Networks (CNNs), due to its capability of learning multi-scale representations and hierarchical features, have been widely used in the field of computer vision. Recent deep learning based edge detectors fully exploit the merits of CNNs and constantly boot the performance of edge detection algorithms in publicly available datasets. In the early stage, the CNN-based edge detectors learned to detect edges from image patches [19, 20], until the proposal of the Holistically-nested edge detection  [21], which used a fully convolutional design and could thus be trained end-to-end and performed image-to-image prediction. Besides, the deep supervision technique used in HED enabled the network to learn and fuse multi-scale and multi-level features, which were important for extracting edges from various scales. Later on, various approaches were proposed following the fashion of HED, but with more complicated techniques to enhance the multi-scale feature learning and fusion cabability of the proposed methods. CEDN [22] trained a fully convolutional encoder-decoder network (the encoder was initialized using the VGG-16 network [23] and followed by a light-weight decoder network), in order to fully exploit the multi-scale representations learned by the pretrained VGG-16 net for object contour detection. Starting from the pretrained VGG net [23] and ResNet [24], COB [25] showed that the boundaries cound be detected more accurately by learning jointly the contour strength and contour orientations in multi-scale. RCF [26] proposed to combine the features extracted from all convolutional layers instead of using only the last layer of each stage as in HED, so as to fully exploit the multi-scale and multi-level information learned by all convolutional layers. By using more training data, the multi-scale testing technique and the improved network architecture, RCF achieved better performance than HED in the publicly available datasets for edge detection. CED [27] proposed a novel refinement network architecture, which increased the resolution of feature maps progressively with sub-pixel convolution, in order to improve the crispness of edges detected by HED. DeepBoundaries [28] extended HED with a multi-resolution architecture, and improved the edge detection accuracy of HED by using more training data and a well designed loss function. The performance of DeepBoundaries was further improved by combining the proposed multi-resolution architecture with the Normalized Cuts technique [29]. Similar to CED, the RefineContourNet (RCN) [30] was also designed following the idea of RefineNet [31], which used a multi-path refinement architecture, for more effective utilization of the high-level abstraction ability of the ResNet. BDCN [32] adopted a bi-directional cascade network structure, which supervised individual layers with edges at its specific scale, to extract edges at different scales. A Scale Enhancement Module (SEM) was also proposed in BDCN in order to enrich the multi-scale representations of each layer by using dilated convolutions. CATS [33] proposed to improve the localization accuracy of existing deep edge detectors with a novel tracing loss and a context-aware fusion block, tackling the problem of feature mixing and side mixing of existing methods.
Despite the efficiency of these end-to-end deep edge detectors, which constantly boosted the performance of edge detection, their performance, however, heavily relies on the pretrained weights of the backbone network on ImageNet dataset [34]. Apparently their performance will drop if the pretrained weights are not used, as much smaller amount of data is used for training. Apart from these methods, DexiNed [35] proposed a new network architecture for edge detection and created a new benchmark dataset (BIPEDv2).It showed that by training all models from scratch in the BIPEDv2 dataset, DexiNed outperformed all existing CNN-based edge detectors. Different from previous methods, DexiNed was not built on the basis of VGG net or the ResNet. Therefore, no pretrained weights on the ImageNet dataset is available for DexiNed, which could be an important reason why it does not provide quantitative comparisons on the previous boundary detection datasets.
As is known to all, the following four aspects are of crucial importance for the performance of a CNN model: the training data, network architecture, loss function and the training strategy (optimizers, hyperparameter settings). The use of pretrained weights on the ImageNet dataset in most existing CNN-based edge detectors makes it difficult to design new network architectures for edge detection, as we have to obtain an optimal pretrained weights for the new network by training it first using the ImageNet dataset, and then fine tune the model using the datasets for edge detection. However, it is quite expensive to train a model on ImageNet dataset, which contains more than one million images, and thus not feasible for many reserchers due to the limited computation resources. Besides, the pretrained weights of the VGG net or the ResNet might not help to improve the edge detection accuracy if the backbone network architecture is modified or our aim is to train the model for edge detection in a new kind of images, Synthetic Aperture Radar (SAR) images for example, because of the strong differences in the statistics between SAR and natural images [9]. Therefore, the use of pretrained weights in existing CNN-based edge detectors limits heavily the design space of CNN models for edge detection. To facilitate the design of new network architectures for edge detection in future work, we set new benchmarks to evaluate edge detection algorithms in publicly available datasets. We train state-of-the-art deep learning based edge detectors in three publicly available datasets for edge detection [36, 37, 35] and compute the optimal performance that can be achieved by them when training from scratch. We then propose a new network architecture for edge detection, the multi-stream and multi-scale fusion net (msmsfnet), and show that it outperforms these edge detectors when training from scratch in the three datasets.
The contributions of this work are two-fold:
@@ I Introduction
  • we set new benchmarks to evaluate edge detection algorithms in publicly available datasets, which facilitates the design of new network architectures for edge detection;
@@
  • we propose a new network architecture for edge detection, and it outperforms state-of-the-art edge detectors in three publicly available datasets when all models are trained from scratch.
@@
The rest of the paper is organized as follows: In Section II, we provide the details about the proposed network architecture; in Section III, we demonstrate the efficiency of the proposed method by comparing it with state-of-the-art deep learning based edge detectors in three publicly available datasets for edge detection. Conclusions are finally given in Section IV.
II Msmsfnet: a multi-stream and multi-scale fusion net for edge detection
II-A Motivation
@@
Developing deep convolutional neural networks for computer vision tasks has become popular ever since the proposal of Alexnet [38], which emphasizes the importance of depth on the performance of CNNs. Alexnet consists of five convolutional layers and three fully connected layers, among which kernels of large sizes ($11\times 1$ 11 and $5\times 5$) are used in the first two convolutional layers. Afterwards, the very deep convolutional networks (VGG net) was proposed in  [23], which utilized small convolutional filters ($3\times 3$) in all convolutional layers. As described in  [23], a sequence of $3\times 3$ convolutional filters has the same effective receptive field as convolutional filters of larger size, while it has fewer parameters to learn when the number of input and output channels is equal. The VGG net [23] showed that the accuracy of the network on the large scale image recognition tasks can be significantly boosted by pushing the depth of the network to 19 weight layers. Parallel to the work of VGG net, GoogLeNet [39] (also known as inceptionv1) achieved similarly high performance to the VGG net but with much more complicated network architecture. GoogLeNet pushed the depth of the network to 22 weight layers through the excessive use of dimension reduction with $1\times 1$ convolutions. Dimension reduction using $1\times 1$ convolution was also widely used in subsequent inception network architectures [40, 41, 42], the ResNet and its variants [43, 44], which reduces the number of parameters to be learned in each layer and thus makes it possible to increase the depth of the network significantly. Besides, the number of parameters in the network can also be reduced by factorizing large convolutional layers into small ones or by factorizing them into spatial asymmetric convolutions [41].
@@
As demonstrated in previous works [38, 23, 39, 24, 40, 41, 42], the performance of CNN models can be boosted significantly by increasing the depth of the network, while a crucial factor to develop very deep CNNs is to reduce the number of parameters in each layer. In this work, we follow the strategy used in [41] and factorize usual convolutional filters into spatial asymmetric convolutions so as to reduce the number of parameters to be learned in each layer. Morespecifically, a $n\times n$ convolutional layer can be factorized into a $1\times n$ convolutional layer followed by a $n\times 1$ convolutional layer (or the reverse). This factorization reduces significantly the number of parameters to be learned while maintaining the size of the receptive field. For instance, factorizing a $3\times 3$ convolution into a $1\times 3$ convolution followed by a $3\times 1$ convolution contributes to a $33$ percent reduction in the number of parameters for the same number of input/output filters (assuming that the number of input and output filters is equal). Much more computational cost can be saved for the factorization of filters with larger kernel sizes. As what we will show in the following, we push the depth of our proposed model to 74 weight layers by utilizing excessively spatial asymmetric convolutions.
II-B Network architecture
For the task of edge detection, the multi-scale representation capability of models is of crucial importance. Existing CNN-based edge detectors constantly boost the performance of edge detection by enriching the multi-scale representations learned by the backbone networks in various ways. Although CNN models are naturally multi-scale/muti-level feature extractors because of the downsampled strided layers as well as the increasingly larger receptive field, we can still improve the multi-scale representation capability of CNNs by exploiting network architectures of higher quality.
@@ II-B Network architecture
To enrich the multi-scale representations, we devise a multi-stream and multi-scale fusion block (msmsfblock), and construct our msmsfnet by simply stacking our proposed msmsfblock. As shown in Figure 1 which displays the configurations of our msmsfblock, the input feature maps are fed into four parallel branches, extracting features at different scales due to the difference in the size of the receptive field. The extracted features are then fused by a two-step fusion stage, where the first step fuses features of closer scales, and the second step produces the final output of the msmsfblock. In our msmsfblock, we utilize excessively spatial asymmetric convolutions ($1\times n$ and $n\times 1$ convolutions), and only the final $3\times 1$ convolutional filter is followed by a ReLU function [38]. Batch normalization [40] is not used in our msmsfblock.

### v3

@@ Abstract
Edge detection is a long-standing problem in computer vision. Despite the efficiency of existing algorithms, their performance, however, rely heavily on the pre-trained weights of the backbone network on the ImageNet dataset. The use of pre-trained weights in previous methods significantly increases the difficulty to design new models for edge detection without relying on existing well-trained ImageNet models, as pre-training the model on the ImageNet dataset is expensive and becomes compulsory to ensure the fairness of comparison. Besides, the pre-training and fine-tuning strategy is not always useful and sometimes even inaccessible. For instance, the pre-trained weights on the ImageNet dataset are unlikely to be helpful for edge detection in Synthetic Aperture Radar (SAR) images due to strong differences in the statistics between optical images and SAR images. Moreover, no dataset has comparable size to the ImageNet dataset for SAR image processing. In this work, we study the performance achievable by state-of-the-art deep learning based edge detectors in publicly available datasets when they are trained from scratch, and devise a new network architecture, the multi-stream and multi-scale fusion net (msmsfnet), for edge detection. We show in our experiments that by training all models from scratch, our model outperforms state-of-the-art edge detectors in three publicly available datasets. We also demonstrate the efficiency of our model for edge detection in SAR images, where no useful pre-trained weight is available. Finally, We show that our model is able to achieve competitive performance on the BSDS500 dataset when the pre-trained weights are used.
keywords:
@@
Edge detection , Deep learning , Pre-trained weights , ImageNet , Msmsfnet , SAR images
journal: Preprint
1 Introduction
Edge detection in both natural images [1, 2, 3, 4, 5, 6] and remote sensing images [7, 8, 9, 10, 11] has been studied for several decades due to the importance of edge features in image analysis and the use of edges in many applications such as line segment detection [12, 13, 14], image segmentation [5, 15], optical flow estimation [16] and object detection [17, 18]. Despite the large amount of research dedicated to the problem of edge detection, state-of-the-art edge detection algorithms mainly rely on deep learning. However, their performance heavily relies on the pre-trained weights of the backbone network on the ImageNet dataset [19]. Apparently their performance will drop if the pre-trained weights are not used, as much smaller amount of data is used for training. As is known to all, the following four aspects are of crucial importance for the performance of deep learning models: the training data, network architecture, loss function and the training strategy (optimizers, hyperparameter settings). The use of pre-trained weights on the ImageNet dataset in most existing edge detectors increases the difficulty to design new network architectures for edge detection without relying on existing well-trained ImageNet models, as it is quite expensive to train a model on the ImageNet dataset (the ImageNet-1k dataset contains around 1.28 million images and the ImageNet-21k dataset contains around 14 million images), and thus not feasible for many researchers due to limited computation resources.
On the other hand, the pre-trained weights of existing ImageNet models (such as VGG net  [20] and ResNet  [21]) are unlikely to improve the edge detection accuracy if the aim is to train the model for edge detection in a different kind of images, Synthetic Aperture Radar (SAR) images for example, because of strong differences in the statistics between SAR and natural images  [9]. Natural images are usually contaminated by weak, additive Gaussian noise. In contrast, SAR images are corrupted by strong multiplicative noise. In a real SAR image, the intensity probability density function (pdf) of a homogeneous area follows a Gamma distribution under the hypothesis of fully developed speckle  [22]. Another strong distinction between a natural image and a real SAR image is the range of pixel values. In a natural image, the maximum pixel value is often 255, while in a real SAR image, the maximum pixel value could be 10000 or much larger. These strong distinctions determine that the pre-trained weights of the backbone networks on the ImageNet dataset are unlikely to help to improve the performance of existing deep learning methods for edge detection in real SAR images. Furthermore, no dataset in the field of SAR image processing has comparable size to the ImageNet dataset. The pre-training (in a very large dataset) and then fine-tuning (in a small dataset) strategy commonly used in the field of computer vision is thus not suitable for SAR image processing. Hence, it’s indeed meaningful to study the performance achievable by deep learning based edge detectors without using pre-trained weights.
To facilitate the design of new network architectures for edge detection in future work, we set new benchmarks to evaluate edge detection algorithms in publicly available datasets. We train state-of-the-art deep learning based edge detectors in three publicly available datasets for edge detection [23, 24, 25] and compute the optimal performance that can be achieved by them when training from scratch. We then propose a new network architecture for edge detection, the multi-stream and multi-scale fusion net (msmsfnet), and show that it outperforms these edge detectors when training from scratch in the three datasets. The efficiency of the proposed method is then demonstrated for edge detection in SAR images, in which case no useful pre-trained weight is available. Finally, we show in our experiments that with the help of pre-trained weights from the ImageNet dataset, the performance of our model on the BSDS500  [23] dataset improves further. This pre-training and fine-tuning experiment provides additional evidence to support our arguments that: 1) the performance of deep learning models rely heavily on the amount of training data; 2) the comparison of the performance of different algorithms should be done under the same experimental settings (i.e., the same amount of data should be used for training) and would be unfair otherwise.
The contributions of this work are summarized as follows:
  • 1.
@@ 1 Introduction
    We set new benchmarks to evaluate edge detection algorithms in publicly available datasets, which facilitates the design of new network architectures for edge detection.
  • 2.
@@
    We propose a new network architecture for edge detection, which outperforms state-of-the-art edge detectors in three publicly available datasets when all models are trained from scratch, and the proposed method is able to achieve competitive performance when the pre-trained weights are used.
  • 3.
    The efficiency of the proposed method is further demonstrated for edge detection in SAR images, where no useful pre-trained weight is available for existing methods. This serves as an important evidence showing that it’s indeed meaningful to study the performance that can be achieved by deep learning based edge detectors when they are trained from scratch.
@@
The rest of the paper is organized as follows: Section 2 gives a brief review of edge detection algorithms for both natural images and SAR images; Section 3 provides the details about the proposed network architecture; in Section 4, we demonstrate the efficiency of the proposed method by comparing it with state-of-the-art deep learning based edge detectors in both natural images and SAR images; Conclusions are finally given in Section 5.
2 Related work
2.1 Edge detection in natural images
Since state-of-the-art methods for edge detection in natural images are mainly based on deep learning, we only give a brief review of the deep learning based edge detection algorithms.
In the early stage, the CNN-based edge detectors learned to detect edges from image patches [26, 27], until the proposal of the Holistically-nested edge detection (HED) [28], which used a fully convolutional design and could thus be trained end-to-end and performed image-to-image prediction. Besides, the deep supervision technique [29] used in HED enabled the network to learn and fuse multi-scale and multi-level features, which were important for extracting edges from various scales. Later on, various approaches were proposed following the fashion of HED, but with more complicated techniques to enhance the multi-scale feature learning and fusion capability of the proposed methods. CEDN [30] trained a fully convolutional encoder-decoder network (the encoder was initialized using the VGG-16 network [20] and followed by a light-weight decoder network), in order to fully exploit the multi-scale representations learned by the pre-trained VGG-16 net for object contour detection. Starting from the pre-trained VGG net [20] and ResNet [21], COB [31] showed that the boundaries could be detected more accurately by learning jointly the contour strength and contour orientations in multi-scale. RCF [32] proposed to combine the features extracted from all convolutional layers instead of using only the last layer of each stage as in HED, so as to fully exploit the multi-scale and multi-level information learned by all convolutional layers. By using more training data, the multi-scale testing technique and the improved network architecture, RCF achieved better performance than HED in the publicly available datasets for edge detection. CED [33] proposed a novel refinement network architecture, which increased the resolution of feature maps progressively with sub-pixel convolution, in order to improve the crispness of edges detected by HED. DeepBoundaries [34] extended HED with a multi-resolution architecture, and improved the edge detection accuracy of HED by using more training data and a well designed loss function. The performance of DeepBoundaries was further improved by combining the proposed multi-resolution architecture with the Normalized Cuts technique [35]. Similar to CED, the RefineContourNet (RCN) [36] was also designed following the idea of RefineNet [37], which used a multi-path refinement architecture, for more effective utilization of the high-level abstraction ability of the ResNet. BDCN [38] adopted a bi-directional cascade network structure, which supervised individual layer with edges at its specific scale, to extract edges at different scales. A Scale Enhancement Module (SEM) was also proposed in BDCN in order to enrich the multi-scale representations of each layer by using dilated convolutions. CATS [39] proposed to improve the localization accuracy of existing deep edge detectors with a novel tracing loss and a context-aware fusion block, tackling the problem of feature mixing and side mixing of existing methods. Apart from these methods, DexiNed [25] proposed a new network architecture for edge detection and created a new benchmark dataset (BIPEDv2). It showed that by training all models from scratch in the BIPEDv2 dataset, DexiNed outperformed all existing CNN-based edge detectors. Different from previous methods, DexiNed was not built on the basis of VGG net or the ResNet. Therefore, no pre-trained weights on the ImageNet dataset is available for DexiNed, which could be an important reason why it does not provide quantitative comparisons on previous boundary detection datasets. UAED  [40] proposed to improve the edge detection accuracy by measuring the uncertainty and investigating the subjectivity and ambiguity among different annotations of the edge ground truth. MUGE  [41] improved further the performance of UAED by estimating the edge granularity of individual annotation, which could be used to guide the generation of diversified edge maps at different controllable granularities. Despite the efficiency of UAED and MUGE, the strategies proposed by them are applicable only when multiple annotations are available for each training image. EDTER  [42] proposed a transformer-based edge detector to fully exploit the long-range dependencies captured by vision transformer  [43], and achieved competitive performance in several benchmark datasets for edge detection. However, the performance achieved by EDTER heavily rely on the pre-trained weights of the backbone network  [43] on the ImageNet-22k dataset, namely the ImageNet-21k dataset plus the ImageNet-1k dataset. Much more data is actually used to train the model than previous methods. As what we will show in the experimental section (Subsection  4.7), the performance of EDTER is far below than that of CNN-based edge detectors when training from scratch.
2.2 Edge detection in SAR images
Edge detection in SAR images has been studied for several decades due to the importance of edge features in image analysis, line segment detection [14], road extraction [44] and many other applications [11]. Although many edge detectors are efficient for optical images [1, 2, 5, 6], they are, however, not suitable for edge detection in SAR images, mainly because of the strong differences in the statistics of optical and SAR images [7]. It has been carefully studied and shown in  [7] that the usual difference based gradient operators, which are commonly used in edge detectors for optical images, do not have a constant false alarm rate (CFAR) for SAR images, while the Ratio of Average (ROA) [7] is more suitable to perform edge detection in SAR images thanks to its CFAR property. Later on, the exponentially weighted average was proved to be optimal to estimate the local mean value of a homogeneous area under the hypothesis of a stochastic multi-edge model, and the Ratio of Exponentially Weighted Average (ROEWA) was proposed to perform edge detection in SAR images [8].  [45] and  [46] were still based on the ratio operation, but use different shapes of window functions to estimate the local mean value of homogeneous areas. Despite the valuable CFAR property of these ratio-based edge detectors, the edge detection results obtained by them were not satisfying enough because of the strong multiplicative noise in SAR images, especially in the most challenging 1-look situation. To improve the edge detection performance in 1-look real SAR images, a deep learning based edge detector GRHED was proposed in  [9]. Due to the absence of a real SAR dataset to train deep learning models for edge detection, it proposed a strategy to enable models trained using simulated SAR images to be efficient for real SAR images.
3 Msmsfnet: a multi-stream and multi-scale fusion net for edge detection
3.1 Motivation
@@
Developing deep convolutional neural networks for computer vision tasks has become popular ever since the proposal of Alexnet [47], which emphasizes the importance of depth on the performance of CNNs. Alexnet consists of five convolutional layers and three fully connected layers, among which kernels of large sizes ($11\times 1$ 11 and $5\times 5$) are used in the first two convolutional layers. Afterwards, the very deep convolutional networks (VGG net) was proposed in  [20], which utilized small convolutional filters ($3\times 3$) in all convolutional layers. As described in  [20], a sequence of $3\times 3$ convolutional filters has the same effective receptive field as convolutional filters of larger size, while it has fewer parameters to learn when the number of input and output channels is equal. The VGG net [20] showed that the accuracy of the network on the large scale image recognition tasks can be significantly boosted by pushing the depth of the network to 19 weight layers. Parallel to the work of VGG net, GoogLeNet [48] (also known as inceptionv1) achieved similarly high performance to the VGG net but with much more complicated network architecture. GoogLeNet pushed the depth of the network to 22 weight layers through the excessive use of dimension reduction with $1\times 1$ convolutions. Dimension reduction using $1\times 1$ convolution was also widely used in subsequent inception network architectures [49, 50, 51], the ResNet and its variants [52, 53], which reduces the number of parameters to be learned in each layer through the use of $1\times 1$ convolution and thus makes it possible to increase the depth of the network significantly. Besides, the number of parameters in the network can also be reduced by factorizing large convolutional layers into small ones or by factorizing them into spatial asymmetric convolutions [50].
@@
As demonstrated in previous work [47, 20, 48, 21, 49, 50, 51], the performance of CNN models can be boosted significantly by increasing the depth of the network, while a crucial factor to develop very deep CNNs is to reduce the number of parameters in each layer. In this work, we follow the strategy used in [50] and factorize usual convolutional filters into spatial asymmetric convolutions so as to reduce the number of parameters to be learned in each layer. More specifically, a $n\times n$ convolutional layer can be factorized into a $1\times n$ convolutional layer followed by a $n\times 1$ convolutional layer (or the reverse). This factorization reduces significantly the number of parameters to be learned while maintaining the size of the receptive field. For instance, factorizing a $3\times 3$ convolution into a $1\times 3$ convolution followed by a $3\times 1$ convolution contributes to a $33$ percent reduction in the number of parameters for the same number of input/output filters (assuming that the number of input and output filters is equal). Much more computational cost can be saved for the factorization of filters with larger kernel sizes. We utilize spatial asymmetric convolutions excessively in our network and push the depth of our proposed model to 74 weight layers.
@@
3.2 Network architecture
For the task of edge detection, the multi-scale representation capability of models is of crucial importance. Existing CNN-based edge detectors constantly boost the performance of edge detection by enriching the multi-scale representations learned by the backbone networks in various ways. Although CNN models are naturally multi-scale/multi-level feature extractors because of the downsampled strided layers as well as the increasingly larger receptive field, we can still improve the multi-scale representation capability of CNNs by exploiting network architectures of higher quality.
To enrich the multi-scale representations, we devise a multi-stream and multi-scale fusion block (msmsfblock), and construct our msmsfnet by simply stacking our proposed msmsfblock. As shown in Figure 1 which displays the configurations of our msmsfblock, the input feature maps are fed into four parallel branches, extracting features at different scales due to the difference in the size of the receptive field. The extracted features are then fused by a two-step fusion stage, where the first step fuses features of closer scales, and the second step produces the final output of the msmsfblock. In our msmsfblock, we utilize spatial asymmetric convolutions excessively ($1\times n$ and $n\times 1$ convolutions), and only the final $3\times 1$ convolutional filter is followed by a ReLU function [47]. Batch normalization [49] is not used in our msmsfblock.